# Exploring Spatial Segregation Induced by Competition Avoidance as Driving Mechanism for Emergent Coexistence in Microbial Communities

### 12 Feb 2024 v2

@@ V Discussion
Similarly, it can be argued that the absence of uneven nutrient distribution in our model may impose limitations on its applicability. Nevertheless, when considering experimental communities, they are typically cultured in agar plates or batch cultures where nutrients are generally thoroughly mixed, resulting in a uniform distribution throughout the medium. Given this context, we assume that there is no substantial nutrient concentration gradient influencing bacterial motion. Thus, we posit that the predominant driving force for movement is the avoidance of competition and we showed how even in isotropic environments with no particular nutrients distribution, is still possible to have a spatial structure with meta-communities, affecting the coexistence among species. A consumer-resource model would necessitate making various assumptions about nutrient abundances, types, temporal dynamics, spatial distribution, and more. The Lotka-Volterra formalism allows the modeling of resource competition without delving into specific resource details. In Chang et al.’s experiments, nutrients start off limited, leading to bacterial competition. However, the nutrients undergo repeated replenishment throughout multiple cycles. If we consider the experiment duration significantly exceeds the duration of a single cycle, it is akin to assuming an overall stable nutrient concentration. Consequently, even if a different model, like MacArthur’s Consumer-Resource model incorporating nutrient dynamics, were employed, assuming globally constant nutrient concentrations over time would essentially reduce it to a formalism equivalent to Lotka-Volterra dynamics.
@@
In summary, our objective with this work is to convey a fundamental message — that space distribution is a crucial factor. Particularly in ecological theoretical models, the common approach involves a mean-field approximation, assuming that each species interacts with all others in the system. While a completely microscopic approach may pose challenges, our study demonstrates that incorporating a mesoscale structure, such as distinct communities or patches in space, can profoundly influence the dynamics and coexistence patterns. Simultaneously, we illustrated how minimal adaptability levels can lead to the formation of isolated bacterial clusters, maintaining a high degree of randomness that mirrors the intricate stimuli influencing bacterial motion. We believe that this work has the potential to pave the way for a new research direction, emphasizing the importance of considering the delicate spatial equilibrium between species within a microbial community as a pivotal element to be incorporated into theoretical models and investigations.

### 26 Feb 2024 v3

@@ V Discussion
Similarly, it can be argued that the absence of uneven nutrient distribution in our model may impose limitations on its applicability. Nevertheless, when considering experimental communities, they are typically cultured in agar plates or batch cultures where nutrients are generally thoroughly mixed, resulting in a uniform distribution throughout the medium. Given this context, we assume that there is no substantial nutrient concentration gradient influencing bacterial motion. Thus, we posit that the predominant driving force for movement is the avoidance of competition and we showed how even in isotropic environments with no particular nutrients distribution, is still possible to have a spatial structure with meta-communities, affecting the coexistence among species. A consumer-resource model would necessitate making various assumptions about nutrient abundances, types, temporal dynamics, spatial distribution, and more. The Lotka-Volterra formalism allows the modeling of resource competition without delving into specific resource details.
@@
In conclusion, we believe that our approach is innovative as it validates the utilization of population-based models integrating meso-scale structures. While a fully microscopic approach pose technical challenges, a mean-field approach would oversimplify, and existing works incorporating meso-structures often do not provide generative mechanisms for them. We introduced a microscopic process that potentially rationalizes the adoption of patch configurations. Simultaneously, we illustrated how minimal adaptability levels can lead to the formation of isolated bacterial clusters, maintaining a high degree of randomness that mirrors the intricate stimuli influencing bacterial motion. We believe that this work has the potential to pave the way for a new research direction, emphasizing the importance of considering the delicate spatial equilibrium between species within a microbial community as a pivotal element to be incorporated into theoretical models and investigations.